OpenStudy (anonymous):

An AC generator supplies an rms voltage of 220 V at 50.0 Hz. It is connected in series with a 0.300 H inductor, a 6.00 μF capacitor and a 346 Ω resistor. What is the impedance of the circuit?

OpenStudy (anonymous):

Please please please help me! :)

OpenStudy (anonymous):

since the load is series connected we can simply add the resistance and given reactance... the total impedance \[Z_{TOTAL}=R+j(X_L-X_C)\] given also is the frequency \(f=50\) Hz, we can determine the inductive reactance \[X_L=2 \pi fL=2 \pi (50)(0.30)\approx94.25 ~Ohms\]and capacitive reactance \[X_C=\frac {1}{2 \pi f C}=\frac{1}{2 \pi (50)(6\times10^{-6})}\approx530.52~Ohms\]therefore, the total impedance of the circuit \[Z_{TOTAL}=346+j(94.25-530.52)=346-j436.27~Ohms\]